A projectile is fired with a speed of 129 m/s at an angle of 49o. Find the vertical component of the initial velocity
1
Expert's answer
2020-03-06T10:27:38-0500

The vertical component of the initial velocity:


"v_{oy}=v_{o}\sin{49}"

"v_{oy}=129\sin{49}=97,4\frac{m}{s}"